# ReactEMG: Zero-Shot, Real-time Intent Detection from sEMG via Masked Modeling

### v5

@@ 4.1 Evaluation metrics
Many prior studies report window-level or segment-level classification accuracy, a metric that does not capture performance in continuous online settings. Approaches that aim to approximate real-time use typically assume unrealistic online conditions and rarely report latency or stability [29]. Online user testing can reveal these effects; however, small sample sizes and substantial human variability make it difficult to control confounding factors and attribute observed differences to the model architecture. To standardize model evaluation in offline for real-time device control, we report three complementary metrics: raw accuracy, transition accuracy, and latency.
@@ 4.2 Baselines
LDA [57] uses a linear discriminant classifier over hand-crafted features computed on sliding windows. The featurization extracts a compact set of discriminative metrics (e.g., RMS, waveform length, and median amplitude spectrum) per channel, standardize them, then apply a linear discriminant projection. Whereas Qi et al. pair their feature transform with a GA-optimized ELM classifier, we instead use the discriminant itself for classification to maintain a simple shallow baseline that produces one label per window.
MLP [49] employs a sliding-window strategy and assigns a single label to each window. The preprocessing stage extracts a compact set of time-domain metrics—such as RMS, variance, mean absolute value, and zero crossings—from each channel in the window. After standardization, these features are fed into a multi-layer perceptron, which outputs one predicted gesture label per window.
LSTM [32] processes EMG as a sequence, using fixed-length segments to train a time-distributed classifier. A fully connected layer first maps the per-timestep inputs to a higher-dimensional representation, which is then fed to an LSTM layer that captures temporal dependencies. A final layer produces gesture class scores at each timestep, yielding dense predictions that are causal within the segment.
ED-TCN [30] is an encoder-decoder temporal convolutional network that relies solely on 1D convolutions. The encoder applies progressively downsampled temporal convolutions with relatively long kernels and temporal max-pooling, while the decoder mirrors this process through upsampling and additional convolutional layers. A final time-distributed dense layer projects the decoder outputs into the target class dimension, producing per-timestep gesture classifications.
TraHGR [33] operates on the raw signal space via patch embeddings. Each window is split into non-overlapping patches along two paths—temporal (per-channel) and featural (across channels)—which are linearly projected to tokens, augmented with learned positional embeddings, and prepended with a class token. The token sequences pass through Transformer encoder blocks; the class-token representations are fused and fed to a linear head to produce one label per window. The model is trained fully supervised with cross-entropy.

### v6

@@ 4.1 Evaluation metrics
Many prior studies report window-level or segment-level classification accuracy, a metric that does not capture performance in continuous online settings. Approaches that aim to approximate real-time use typically assume unrealistic online conditions and rarely report latency or stability [BaronaLpez2024]. Online user testing can reveal these effects; however, small sample sizes and substantial human variability make it difficult to control confounding factors and attribute observed differences to the model architecture. To standardize model evaluation in offline for real-time device control, we report three complementary metrics: raw accuracy, transition accuracy, and latency.
@@ 4.2 Baselines
LDA [Qi2019-xt] uses a linear discriminant classifier over hand-crafted features computed on sliding windows. The featurization extracts a compact set of discriminative metrics (e.g., RMS, waveform length, and median amplitude spectrum) per channel, standardize them, then apply a linear discriminant projection. Whereas Qi et al. pair their feature transform with a GA-optimized ELM classifier, we instead use the discriminant itself for classification to maintain a simple shallow baseline that produces one label per window.
MLP [Lee2021] employs a sliding-window strategy and assigns a single label to each window. The preprocessing stage extracts a compact set of time-domain metrics—such as RMS, variance, mean absolute value, and zero crossings—from each channel in the window. After standardization, these features are fed into a multi-layer perceptron, which outputs one predicted gesture label per window.
LSTM [Simo2019] processes EMG as a sequence, using fixed-length segments to train a time-distributed classifier. A fully connected layer first maps the per-timestep inputs to a higher-dimensional representation, which is then fed to an LSTM layer that captures temporal dependencies. A final layer produces gesture class scores at each timestep, yielding dense predictions that are causal within the segment.
ED-TCN [Betthauser2020] is an encoder-decoder temporal convolutional network that relies solely on 1D convolutions. The encoder applies progressively downsampled temporal convolutions with relatively long kernels and temporal max-pooling, while the decoder mirrors this process through upsampling and additional convolutional layers. A final time-distributed dense layer projects the decoder outputs into the target class dimension, producing per-timestep gesture classifications.
TraHGR [Zabihi2023] operates on the raw signal space via patch embeddings. Each window is split into non-overlapping patches along two paths—temporal (per-channel) and featural (across channels)—which are linearly projected to tokens, augmented with learned positional embeddings, and prepended with a class token. The token sequences pass through Transformer encoder blocks; the class-token representations are fused and fed to a linear head to produce one label per window. The model is trained fully supervised with cross-entropy.